# On tori periods of Weil representations of unitary groups

### v1

@@ 4.1.1. Splitting of the embedding $\iota$
The goal of this subsection is to describe the data needed in order to construct a splitting $\tilde{\iota}\colon\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)\times\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)\to\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ of $\iota$. Let $\mathbb{A}_{\mathbf{K}}$ be the adeles of $\mathbf{K}$, and let $\omega_{\mathbf{K}/\mathbf{F}}$ be the quadratic character attached to the field extension $\mathbf{K}/\mathbf{F}$ by global class field theory. Similarly to Section 3.1.1, in order to construct a splitting, we need to choose automorphic characters $\chi_{\mathrm{V}}=\bigotimes_{v}\chi_{\mathrm{V}_{v}}$ and $\chi_{\mathrm{W}}=\bigotimes_{v}\chi_{\mathrm{W}_{v}}$ of $\mathbf{K}^{\times}\backslash\mathbb{A}_{\mathbf{K}}^{\times}$, such that
@@ 4.1.2. Theta lifting for unitary groups of $1$-dimensional spaces over étale algebras
Similarly to Section 3.1.2, we define a theta correspondence for unitary groups of one-dimensional spaces over an étale algebra. We use the definitions in Section 2.2 with $F=\mathbf{F}$ and $K=\mathbf{K}$. Let $\delta\in\mathbf{K}^{\times}$ be a trace zero element. If $\mathbf{E}/\mathbf{F}$ is a finite field extension, then $\mathbb{A}_{\mathbf{E}}=\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}$. Recall that in this case, if $\lambda\in\mathbf{E}^{\times}$ and if $R$ is a ring over $\mathbf{F}$, then
@@
In particular, we have that the $\mathbf{F}$-adelic points of the unitary group $\operatorname{U}(L_{\mathbf{E},\lambda})$ are the same as the $\mathbf{E}$-adelic points of $\operatorname{U}_{\mathbf{E}}(L_{\mathbf{E},\lambda})$, the unitary group of $L_{\mathbf{E},\lambda}$ defined over $\mathbf{E}$, i.e.,
@@
Let $\mathbf{E}/\mathbf{F}$ be an étale algebra of rank $n$ over $\mathbf{F}$. As before, we write $\mathbf{E}=\prod_{j=1}^{m}\mathbf{F}_{j}$, where for every $j$, $\mathbf{F}_{j}/\mathbf{F}$ is a finite field extension. We will assume that $L_{\mathbf{E}}^{1}$ is anisotropic, i.e., we will assume that $L_{\mathbf{F}_{j}}$ is a field for every $j$. This is equivalent to the assumption that there is no embedding of $\mathbf{F}$-algebras $\mathbf{K}\hookrightarrow\mathbf{E}$. As in Section 3.1.2, for $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in\mathbf{E}^{\times}$ and $\lambda^{\prime}=\left(\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{m}\right)\in\mathbf{E}^{\times}$, we have that
@@
For every $1\leq j\leq m$, let $\mathrm{V}_{j}=L_{\mathbf{F}_{j},\lambda_{j}}$ and $\mathrm{W}_{j}=\mathrm{W}_{\mathbf{F}_{j},\lambda^{\prime}_{j}}^{\delta}$. Every automorphic character $\alpha\colon\left[\operatorname{U}(L_{\mathbf{E},\lambda})\right]\to\mathbb{C}^{\times}$ is equivalent to a choice $\left(\alpha_{1},\dots,\alpha_{m}\right)$, where $\alpha_{j}\colon\left[\operatorname{U}(\mathrm{V}_{j})\right]\rightarrow\mathbb{C}^{\times}$ is an automorphic character for $1\leq j\leq m$. We will use the usual global theta correspondence to define a theta correspondence for the adelic groups $\operatorname{U}(L_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ and $\operatorname{U}(\mathrm{W}_{\mathbf{E},\lambda^{\prime}}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$. Let $\chi_{L_{\mathbf{E},\lambda}},\chi_{\mathrm{W}_{\mathbf{E},\lambda^{\prime}}^{\delta}}\colon L_{\mathbf{E}}^{\times}\backslash\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\mathbb{C}^{\times}$ be automorphic characters such that
@@
Here $\psi_{j}\colon\mathbf{F}_{j}\backslash\mathbb{A}_{\mathbf{F}_{j}}\to\mathbb{C}^{\times}$ is the character $\psi_{j}=\psi\circ\operatorname{tr}_{\mathbf{F}_{j}/\mathbf{F}}$, where $\operatorname{tr}_{\mathbf{F}_{j}/\mathbf{F}}\colon\mathbf{F}_{j}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\to\mathbb{A}_{\mathbf{F}}$ is the trace map. Denote $\tilde{\iota}=\left(\tilde{\iota}_{1},\dots,\tilde{\iota}_{m}\right)$. We define the global theta lift of $\alpha$ as above by the formula
@@ 4.3. Global theta lifts for unitary groups of one-dimensional spaces
In this section, we discuss the definition of the central value of an $L$-function associated with an automorphic character of $\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}$. Assume first that $\mathbf{E}/\mathbf{F}$ is a field extension. Then $L_{\mathbf{E}}=\mathbf{E}\otimes_{\mathbf{F}}\mathbf{K}$ is a field extension (as we assume that $L_{\mathbf{E}}^{1}$ is anisotropic). We have that $L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}=\mathbb{A}_{L_{\mathbf{E}}}$, and therefore an automorphic character of $\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}$ is the same as an automorphic character of $\mathbb{A}_{L_{\mathbf{E}}}^{\times}$. For an automorphic character $\chi\colon L_{\mathbf{E}}^{\times}\backslash{\mathbb{A}_{L_{\mathbf{E}}}^{\times}}\to\mathbb{C}^{\times}$, we define
@@
4.3.2. Base change for characters of $L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$
@@
We have that
$L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)={\prod_{v}}^{\prime}L_{\mathbf{E}}^{1}\left(\mathbf{F}_{v}\right)={\prod_{v}}^{\prime}\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}\right)^{1}.$
@@
$j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\colon\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}/\left(\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{1}=L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$
given by $j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\left(x\right)=\frac{x}{x^{c}}$ decomposes as $j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}=\bigotimes_{v}j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}$, and therefore is an isomorphism. Analogously to Section 3.3.2, given an automorphic character $\beta\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$, we define an automorphic character $\beta_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\colon L_{\mathbf{E}}^{\times}\backslash\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\mathbb{C}^{\times}$ by the formula
@@ 4.3.3. Non-vanishing of global theta lifts
In this section, we recall a result regarding the non-vanishing of the global theta lift. This result serves as an analog of Theorem 3.4. It is established using the Rallis inner product formula. Since we do not need the generality of the Rallis inner product formula, we will just state the non-vanishing result in the generality we need. Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate one-dimensional hermitian and skew-hermitian spaces over $\mathbf{K}$, respectively. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{\mathbf{F}}^{1}$ and $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(\mathrm{V})\to L_{\mathbf{F}}^{1}$ be admissible isomorphisms. Let $\alpha\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ be an automorphic character. The following theorem due to Yamana follows from [34, Lemma 10.2] [^2] We warn the reader that in [34], $G=\operatorname{U}(\mathrm{W})$ and $H=\operatorname{U}(\mathrm{V})$ in our notation..

### v2

@@ 4.1.1. Splitting of the embedding $\iota$
The goal of this subsection is to describe the data needed in order to construct a splitting $\tilde{\iota}\colon\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)\times\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)\to\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ of $\iota$. Let $\mathbb{A}_{\mathbf{K}}$ be the adeles of $\mathbf{K}$, and let $\omega_{\mathbf{K}/\mathbf{F}}$ be the quadratic character attached to the field extension $\mathbf{K}/\mathbf{F}$ by global class field theory.
Similarly to Section 3.1.1, in order to construct a splitting, we need to choose automorphic characters $\chi_{\mathrm{V}}=\bigotimes_{v}\chi_{\mathrm{V}_{v}}$ and $\chi_{\mathrm{W}}=\bigotimes_{v}\chi_{\mathrm{W}_{v}}$ of $\mathbf{K}^{\times}\backslash\mathbb{A}_{\mathbf{K}}^{\times}$, such that
@@
4.1.2. Notation for theta lifts of automorphic characters of $\left[L_{\mathbf{F}}^{1}\right]$
We introduce the global counterpart of the notation in Section 3.1.2.
Let $\delta\in\mathbf{K}^{\times}$ be a trace zero element, and let $\mu\colon\mathbf{K}^{\times}\backslash\mathbb{A}_{\mathbf{K}}^{\times}\to\mathbb{C}^{\times}$ be an automorphic character such that $\mu_{\restriction_{\mathbb{A}_{\mathbf{F}}^{\times}}}=\omega_{\mathbf{K}/\mathbf{F}}$. Suppose that $\mathrm{V}$ is a hermitian space over $\mathbf{K}$ and that $\beta\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ is an automorphic character (see Section 2.5). We denote
$\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)=\Theta_{\mathrm{W}_{\mathbf{F},1}^{\delta},\mathrm{V},\tilde{\iota}_{\mu},\psi}\left(\beta\circ i^{\prime}_{\mathrm{W}_{\mathbf{F},1}^{\delta}}\right),$
where $i^{\prime}_{\mathrm{W}_{\mathbf{F},1}^{\delta}}\colon\operatorname{U}(\mathrm{W}_{\mathbf{F},1}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)\to L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$ is the obvious isomorphism and where $\tilde{\iota}_{\mu}$ is the splitting associated to the characters $\left(\mu,\mu^{\dim V}\right)$.
@@ 4.1.3. Theta lifting for unitary groups of $1$-dimensional spaces over étale algebras
Similarly to Section 3.1.3, we define a theta correspondence for unitary groups of one-dimensional spaces over an étale algebra.
We use the definitions in Section 2.2 with $F=\mathbf{F}$ and $K=\mathbf{K}$. Let $\delta\in\mathbf{K}^{\times}$ be a trace zero element.
If $\mathbf{E}/\mathbf{F}$ is a finite field extension, then $\mathbb{A}_{\mathbf{E}}=\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}$. Recall that in this case, if $\lambda\in\mathbf{E}^{\times}$ and if $R$ is a ring over $\mathbf{F}$, then
@@
In particular, we have that the $\mathbf{F}$-adelic points of the unitary group $\operatorname{U}(L_{\mathbf{E},\lambda})$ are the same as the $\mathbf{E}$-adelic points of the unitary group of $L_{\mathbf{E},\lambda}$ defined over $\mathbf{E}$, which we denote $\operatorname{U}_{\mathbf{E}}(L_{\mathbf{E},\lambda})$, i.e.,
@@
This relation will allow us to make use of statements about theta lifts of characters of $\operatorname{U}_{\mathbf{E}}(L_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{E}}\right)$.
Let $\mathbf{E}/\mathbf{F}$ be an étale algebra of rank $n$ over $\mathbf{F}$. As before, we write $\mathbf{E}=\prod_{j=1}^{m}\mathbf{F}_{j}$, where for every $j$, $\mathbf{F}_{j}/\mathbf{F}$ is a finite field extension. We will assume that $L_{\mathbf{E}}^{1}$ is anisotropic, i.e., we will assume that $L_{\mathbf{F}_{j}}$ is a field for every $j$. This is equivalent to the assumption that there is no embedding of $\mathbf{F}$-algebras $\mathbf{K}\hookrightarrow\mathbf{E}$.
As in Section 3.1.3, for $\lambda=\left(\lambda_{1},\dots,\lambda_{m}\right)\in\mathbf{E}^{\times}$ and $\lambda^{\prime}=\left(\lambda^{\prime}_{1},\dots,\lambda^{\prime}_{m}\right)\in\mathbf{E}^{\times}$, we have that
@@
For every $1\leq j\leq m$, let $\mathrm{V}_{j}=L_{\mathbf{F}_{j},\lambda_{j}}$ and $\mathrm{W}_{j}=\mathrm{W}_{\mathbf{F}_{j},\lambda^{\prime}_{j}}^{\delta}$. Every automorphic character $\alpha\colon\left[\operatorname{U}(L_{\mathbf{E},\lambda})\right]\to\mathbb{C}^{\times}$ is equivalent to a choice $\left(\alpha_{1},\dots,\alpha_{m}\right)$, where $\alpha_{j}\colon\left[\operatorname{U}(\mathrm{V}_{j})\right]\to\mathbb{C}^{\times}$ is an automorphic character for $1\leq j\leq m$. We will use the usual global theta correspondence to define a theta correspondence for the adelic groups $\operatorname{U}(L_{\mathbf{E},\lambda})\left(\mathbb{A}_{\mathbf{F}}\right)$ and $\operatorname{U}(\mathrm{W}_{\mathbf{E},\lambda^{\prime}}^{\delta})\left(\mathbb{A}_{\mathbf{F}}\right)$.
Let $\chi_{L_{\mathbf{E},\lambda}},\chi_{\mathrm{W}_{\mathbf{E},\lambda^{\prime}}^{\delta}}\colon L_{\mathbf{E}}^{\times}\backslash\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\mathbb{C}^{\times}$ be automorphic characters such that
@@
Here $\psi_{j}\colon\mathbf{F}_{j}\backslash\mathbb{A}_{\mathbf{F}_{j}}\to\mathbb{C}^{\times}$ is the character $\psi_{j}=\psi\circ\operatorname{tr}_{\mathbf{F}_{j}/\mathbf{F}}$, where $\operatorname{tr}_{\mathbf{F}_{j}/\mathbf{F}}\colon\mathbf{F}_{j}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\to\mathbb{A}_{\mathbf{F}}$ is the trace map. Denote $\tilde{\iota}=\left(\tilde{\iota}_{1},\dots,\tilde{\iota}_{m}\right)$.
We define the global theta lift of $\alpha$ as above by the formula
@@
Remark 4.1.
Similarly to Remark 3.4, we have that the kernel of the map $\prod_{j=1}^{m}\operatorname{Mp}_{\psi_{j}}\left(\mathrm{V}_{j},\mathrm{W}_{j}\right)\left(\mathbb{A}_{\mathbf{F}_{j}}\right)\to\operatorname{Mp}_{\psi}\left(\mathrm{V}_{\mathbf{E},\lambda},\mathrm{W}_{\mathbf{F},1}^{\delta}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$ consists of tuples $\left(g_{1},\dots,g_{m}\right)$ such that for every $j$, the projection of $g_{j}$ to $\operatorname{Sp}\left(\operatorname{Res}_{L_{\mathbf{F}_{j}}/\mathbf{F}_{j}}\left(\mathrm{V}_{j}\otimes_{L_{\mathbf{F}_{j}}}\mathrm{W}_{j}\right)\right)$ is the identity and such that if $t_{j}$ is the projection of $g_{j}$ to $\mathbb{S}^{1}$ then $\prod_{j=1}^{m}t_{j}=1$.
@@ 4.3. Global theta lifts for unitary groups of one-dimensional spaces
In this section, we discuss the definition of the central value of an $L$-function associated with an automorphic character of $\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}$.
Assume first that $\mathbf{E}/\mathbf{F}$ is a field extension. Then $L_{\mathbf{E}}=\mathbf{E}\otimes_{\mathbf{F}}\mathbf{K}$ is a field extension (as we assume that $L_{\mathbf{E}}^{1}$ is anisotropic). We have that $L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}=\mathbb{A}_{L_{\mathbf{E}}}$, and therefore an automorphic character of $\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}$ is the same as an automorphic character of $\mathbb{A}_{L_{\mathbf{E}}}^{\times}$. For an automorphic character $\chi\colon L_{\mathbf{E}}^{\times}\backslash{\mathbb{A}_{L_{\mathbf{E}}}^{\times}}\to\mathbb{C}^{\times}$, we define
@@
4.3.2. Base change for characters of $\operatorname{Res}_{\mathbf{E}/\mathbf{F}}L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$
@@
Recall the definition of $\operatorname{Res}_{\mathbf{E}/\mathbf{F}}L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$ from Section 2.5. We have that
$\operatorname{Res}_{\mathbf{E}/\mathbf{F}}L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)={\prod_{v}}^{\prime}\operatorname{Res}_{\mathbf{E}/\mathbf{F}}L_{\mathbf{E}}^{1}\left(\mathbf{F}_{v}\right)={\prod_{v}}^{\prime}\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}\right)^{1}.$
@@
$j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\colon\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}/\left(\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{1}=\operatorname{Res}_{\mathbf{E}/\mathbf{F}}L_{\mathbf{E}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$
given by $j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\left(x\right)=\frac{x}{x^{c}}$ decomposes as $j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}=\bigotimes_{v}j_{\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}$, and therefore is an isomorphism.
Analogously to Section 3.3.2, given an automorphic character $\beta\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$, we define an automorphic character $\beta_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\colon L_{\mathbf{E}}^{\times}\backslash\left(L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}\right)^{\times}\to\mathbb{C}^{\times}$ by the formula
@@ 4.3.3. Non-vanishing of global theta lifts
In this section, we recall a result regarding the non-vanishing of the global theta lift. This result serves as an analog of Theorem 3.5. It is established using the Rallis inner product formula. Since we do not need the generality of the Rallis inner product formula, we will just state the non-vanishing result in the generality we need.
Let $\mathrm{V}$ and $\mathrm{W}$ be non-degenerate one-dimensional hermitian and skew-hermitian spaces over $\mathbf{K}$, respectively. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{\mathbf{F}}^{1}$ and $i^{\prime}_{\mathrm{V}}\colon\operatorname{U}(\mathrm{V})\to L_{\mathbf{F}}^{1}$ be the obvious isomorphisms. Let $\alpha\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ be an automorphic character. The following theorem due to Yamana follows from [35, Lemma 10.2] [^2] We warn the reader that in [35], $G=\operatorname{U}(\mathrm{W})$ and $H=\operatorname{U}(\mathrm{V})$ in our notation..